4. EcoSacks manufactures cloth shopping bags. The controller is preparing a budget for the coming year...

4.

EcoSacks manufactures cloth shopping bags. The controller is preparing a budget for the coming year and asks for your assistance. The following costs and other data apply to bag production.

Direct materials per bag
1.50 yard cotton at $4.50 per yard
0.70 yards canvas finish at $12.50 per yard
Direct labor per bag
1.00 hour at $18.50 per hour
Overhead per bag
Indirect labor $ 1.10
Indirect materials 0.70
Power 0.90
Equipment costs 1.80
Building occupancy 1.40
Total overhead per unit $ 5.90

You learn that equipment costs and building occupancy are fixed and are based on a normal production of 650,000 units per year. Other overhead costs are variable. Plant capacity is sufficient to produce 850,000 units per year.

Labor costs per hour are not expected to change during the year. However, the cotton supplier has informed EcoSacks that it will impose a 20 percent price increase at the start of the coming budget period. No other costs are expected to change.

During the coming budget period, EcoSacks expects to sell 590,000 bags. Finished goods inventory is targeted to increase from the current balance of 170,000 units to 260,000 units to prepare for an expected sales increase the year after next as a result of legislation in several states regarding plastic bags. Production will occur evenly throughout the year. Inventory levels for cotton and canvas are expected to remain unchanged throughout the year. There is no work-in-process inventory.

Required:

a. Prepare a production budget for the coming year.
b. Estimate the materials, labor, and overhead costs for the coming year.
Answer #1
Production Budget
Beginning Stock                                               170,000
Expected Sales                                               590,000
Closing stock                                             (260,000)
Budgeted Production                                               500,000
Part-b
Cost Elements Quantity Per Unit Cost Amount
Direct Materials
Cotton (1.5*500000)=750000 ($4.50*1.20)=$5.4 $405,000.00
Canvas Finish 0.70*500000=350000 $12.50 $4,375,000.00
Total Material Cost $4,780,000.00
Direct labour (500000*1 Hour)=500000 $18.50 $9,250,000.00
Total Laboour Cost $9,250,000.00
Overhead Cost
Indirect labour 500000 $1.10 $550,000.00
Indirect Materials 500000 $0.70 $350,000.00
Power 500000 $0.90 $450,000.00
Equipment Cost 650000 $1.80 $1,170,000.00
Building occupancy 650000 $1.40 $910,000.00
Total Overhead Cost $3,430,000.00
Total production Cost $17,460,000.00